Q22-A sample of fluorine gas has a volume of 5.28 L at 13°C and standard pressure. What will be the pressure in atm if the moles of gas and the temperature do not change but the volume is expanded to 9.82 L? answer in atm.
Answer #1

Standard pressure is taken as 1 atm = P1

Initial volume is = V1 = 5.28 L

Keeping temperatures and moles fixed,  

Final volume = V2 = 9.82 L

Let final pressure = P2

P2/P1 = V1/V2

P2 = (5.28/9.82) ×1 = 0.537 atm

Final pressure = 0.537 atm